# Semantically-Shifted Incremental Adapter-Tuning is A Continual ViTransformer

### v1

@@ 4.3 Ablation Study
Baselines with different PET methods: Tab. 4 shows the results of baselines with three different parameter-efficient tuning methods in each incremental session. It can be observed that the pre-trained model with an adapter achieves the best performance in terms of both the last session accuracy and average accuracy. Fig. 1 demonstrates that tuning with an adapter achieves a better balance between learning new classes and retaining knowledge of old classes. Both VPT-deep and SSF methods tend to prioritize learning new categories, which leads to increased forgetting of previously learned categories. Although VPT-shallow performs well on CIFAR, its limited parameters hinder the model from incrementally learning new classes on ImageNetR. More results on the other datasets can be found in the Supp.
@@
Unified classifier retraining vs. Separate local classifier: As we train separate task-specific classifiers in each incremental session, we propose to retrain the classifier to find the optimal decision boundary for all the classes. Tab. 5 displays the ablation experiments of the classifier re-trained on ImageNetA which is the most difficult benchmark. It can be observed that whether it is a linear or a cosine classifier, retraining the classifier leads to a significant performance improvement. Additionally, incorporating the computation of prototype semantic shifts further enhances the performance by an additional $2\%$ in the cosine classifier. Compared to the classifier alignment methods that do not involve computing updated prototypes, our method demonstrates its superiority as the incremental stages progress. More results on the other datasets can be found in the Supp.
@@
Progressively tuning vs. first session adaptation: Tab. 6 shows the linear probing results of different adaption ways. After finishing the training of the last session, we freeze the pre-trained backbone and only train the classifier using all the samples. It is evident that not performing tuning and solely freezing the pre-trained model leads to the worst performance, regardless of the dataset. First-session adaptation proves to be a good choice as it reduces training time and works well for datasets like CIFAR100 and CUB200. However, for datasets such as ImageNetA and ImageNetR, which have significant domain gaps from the pre-trained model, relying solely on first-session adaptation is suboptimal. By continuously fine-tuning the adapter, we observe that the backbone exhibits stronger representability compared to only tuning in the first session.
@@
Different structures of the adapter: In this paper, we follow AdaptFormer [4] to use parallel adapterMLP as the adapter structure. We also delve deeper into different adapter structures such as Adapter [13] and Convpass [16]. Although these different tuning structures may exhibit performance differences under static settings, the performance differences among those adapter structures are minimal in the context of CIL shown in Tab. 7. This offers us the flexibility to employ various adapter structures within the context of the CIL paradigm.
Comparison to traditional CIL methods: We conduct evaluations by comparing our approach to SOTA traditional CIL methods shown in Tab. 8. We replace the Resnet backbone with the pre-trained ViT model for fair comparison. The results indicate that the performance of iCaRL tends to be inferior compared to SOTA model expansion methods and our proposed method, even when past samples are stored. It can be observed that methods such as Foster and Der, which dynamically expand feature extraction networks, achieve impressive results on ImageNetR. The average accuracy of these methods is only $2\%$ lower than our method. However, on ImageNetA, where there are few-shot samples for many classes, these methods exhibit low performance. More ablation experiments related to hyper-parameters can be found in the supp.
@@
SSIAT (Ours) $79.38_{\pm 0.59}$ $83.63_{\pm 0.43}$ $62.43_{\pm 1.63}$ $70.83_{\pm 1.63}$
Table 8: Comparison to traditional CIL methods on ImageNetR and ImageNetA dataset.
5 Conclusion
Class-incremental learning on a pre-trained model has received significant attention in recent years. In this paper, we first revisit different PET methods in the context of CIL. Then, we propose that incrementally tuning the shared adapter and local classifier without constraints exhibits less forgetting and gains plasticity for learning new classes. Moreover, to train a unified classifier, we calculate the semantic shift of old prototypes and retrain the classifier using updated prototypes in each session. The proposed method eliminates the need for constructing an adapter pool and avoids retaining any image samples. Experimental results on five benchmarks demonstrate the effectiveness of our method which achieves the SOTA performance.
Acknowledgement. This research was supported by Natural Science Fund of Hubei Province (Grant # 2022CFB823), Alibaba Innovation Research program under Grant Contract # CRAQ7WHZ11220001-20978282, and HUST Independent Innovation Research Fund (Grant # 2021XXJS096).
Appendix
A. More experiments on parameter sensitivity
Figure 6: left: Parameter sensitivity of the different modules of task 2 and task 3. right: Incremental performance of varying degrees of update limits on important parameters.
We conduct parameter sensitivity experiments on ImageNetR, and the results are shown in Fig. 6. The left graph illustrates the parameter sensitivity of the second and third tasks. Similar to the trends observed in task one and two as shown in the main paper, the parameter sensitivity of different modules in task two and three also exhibits a high degree of similarity. Therefore, restricting parameter updates based on parameter sensitivity negatively impacts the learning of new categories. To investigate this phenomenon, we conduct experiments, as shown in the right graph. It can be seen that as the restriction on parameter updates in increases, the overall performance decreases.
B. More Ablation Experiments
Adapter dimension and layers insert: As shown in Tab. 9 and 10 left, we further conduct experiments on the specific position to insert the adapter module on the ImageNetR and ImageNetA datasets. We observe that the performance is progressively improved with the number of layers increased. Thus, we insert adapter modules in all $12$ layers in our comparative experiments of various methods. We also conduct ablation experiments on the middle dimension in the adapter in Tab. 9 and 10 right. We observe that increasing the dimension has a positive effect on the performance of the model. Interestingly, setting the middle dimension to 32 did not result in a significant decrease in performance. On the other hand, setting it to 256 led to an improvement in performance but also quadrupled the number of parameters. To strike a balance between performance and the number of fine-tuning parameters, we set the middle dimension to $64$.
@@ B. More Ablation Experiments
Analysis of margin and scale: We conduct experiments on the hyper-parameter of the scale and margin in our cosine loss as shown in Tab. 13. We discover that appropriately increasing scale can enhance the performance of the model. For example, on the ImageNetR dataset, when the scale is set to $20$, the average accuracy is $3.36\%$ higher than when the value is $10$. We also conduct experiments to analyze the influence of the margin. As shown in Tab. 14, we observe that different datasets have different appropriate margin values. For example, on the ImageNetR dataset, we set the margin to $0.0$, and on the CUB200 dataset, we set it to $0.1$.
Layers Form Acc Num #Param Acc
1-3 parallel 80.53 32 0.60M 81.98
1-6 parallel 81.63 64 1.19M 82.09
1-12 parallel 81.95 256 4.87M 82.38
Table 9: Experimental results of the inserted layers of the adapter and the middle dimension.
Different PET methods in the CIL: The experiment results on CIFAR100 are in the main paper. In the supplementary material, we provide experiments with different PET methods on ImageNetR as shown in Tab. 15. We can observe that in the initial sessions, the performance of SSF surpasses that of the adapter. However, due to the tendency of SSF to overfit to the current session classes, there is a significant decline in subsequent incremental sessions and the adapter performs best in both the accuracy of the last session and average accuracy. We also provide experiments on ImageNetA as shown in Tab. 16. We can draw the same conclusion from ImageNetR.
Layers Form Acc Num #Param Acc
1-3 parallel 65.25 32 0.60M 66.49
1-6 parallel 65.81 64 1.19M 66.85
1-12 parallel 66.67 256 4.87M 67.54
Table 10: Experimental results of the inserted layers of the adapter and the middle dimension.
Unified classifier retraining vs. Separate local classifier: The experiment results on ImageNetA are in the main paper. Here we show the accuracy of each session of three different seeds on ImageNetA as shown in Tab. 17. It can be seen that retraining the classifier can improve the performance by $2\%$ to $3\%$ on three seeds, effectively improving the performance of the classifier. Furthermore, classifier retraining with semantic shift estimation can further improve performance by $2\%$ to $3\%$. We also show the results on CUB200 as shown in Tab. 18. The same trend is shown in this dataset. CA can significantly improve the performance, and SSCA can align the prototype and further improve performance. The results on ImageNetR are shown in Tab. 19.
Different pre-trained models. We experiment with pre-trained models (PTMs) with different generalization abilities on ImageNetR and ImageNetA datasets shown in Tab. 11. It can be observed that our method generalizes well to various PTMs. The large-based ViT model can get better performance.
Different tuning methods with SSCA. We incorporate classifier alignment with semantic shift estimation into SSF and prompt tuning shown in Tab. 12. It can be seen that both the performance of prompt-based and SSF tuning approaches show significant improvement. However, our proposed method still outperforms them by a large margin. The results further verify the effectiveness of our proposed method.
Pre-trained Model ImageNetR ImageNetA
ViT-base 1K $80.15_{\pm 0.41}$ $83.87_{\pm 0.26}$ $64.88_{\pm 1.11}$ $72.68_{\pm 1.72}$
ViT-base 21k $79.38_{\pm 0.59}$ $83.63_{\pm 0.43}$ $62.43_{\pm 1.63}$ $70.83_{\pm 1.63}$
ViT-large 21k $83.62_{\pm 0.41}$ $86.70_{\pm 0.69}$ $68.38_{\pm 2.25}$ $74.85_{\pm 1.93}$
Table 11: Results on different pre-trained models on ImageNetR/A.
@@
Table 12: Results for different PET methods on ImageNetR/A.
Scale ImageNetR ImageNetA CIFAR100 CUB200
Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$ Last $\uparrow$ Avg $\uparrow$
s=10 73.80 80.84 56.95 66.88 89.49 93.61 86.39 91.61
s=15 77.83 83.36 59.84 69.25 91.02 94.53 88.46 92.57
s=20 79.55 84.20 60.76 69.62 91.62 94.75 88.51 92.83
s=30 78.90 83.30 60.50 68.73 91.21 94.51 88.60 92.64
Table 13: Experimental results of the influence of scale in cosine loss on different datasets.
Margin ImageNetR ImageNetA CIFAR100 CUB200
[…]
m = 0 79.55 84.20 60.76 69.62 91.62 94.75 88.13 92.33
m = 0.1 78.38 84.25 63.00 72.13 91.60 94.71 88.46 92.57
m = 0.2 76.48 82.94 62.74 73.18 89.71 93.28 88.38 92.56
m = 0.3 73.90 80.76 62.61 72.46 86.83 91.45 87.57 91.89
Table 14: Experimental results of the influence of margin in cosine loss on different datasets.
[…]
SSF [22] 0.2M 94.75 89.28 86.11 82.61 80.19 78.67 77.27 75.4 75.04 72.78 81.21
VPT-deep [15] 0.046M 87.23 64.19 59.16 40.27 39.4 36.6 31.51 33.32 33.79 31.62 45.71
VPT-shallow [15] 0.004M 81.86 73.20 68.81 66.85 64.89 63.81 62.84 62.21 61.35 58.97 66.48
Adapter [4] 1.19M 91.87 88.42 86.51 84.43 82.75 81.51 80.99 80.62 79.75 78.28 83.51
Table 15: Experimental results for baselines with different parameter efficient tuning methods on ImageNetR. We report the overall performance of each session and the average performance.
[…]
SSF [22] 0.2M 82.86 76.11 67.44 65.62 61.47 58.97 54.33 52.32 51.04 51.28 62.14
VPT-deep [15] 0.046M 61.14 31.67 15.55 15.62 10.11 6.05 4.82 4.17 3.37 3.42 15.92
VPT-shallow [15] 0.004M 80.00 70.00 64.92 60.32 56.84 54.26 52.63 51.52 49.68 48.26 58.84
Adapter 1.19M 82.86 74.72 71.43 67.62 64.68 62.15 59.14 56.89 55.20 55.50 65.02
Table 16: Experimental results for baselines with different parameter efficient tuning methods on ImageNetA. We report the overall performance of each session and the average performance.
Seed Method Ses.1 Ses.2 Ses.3 Ses.4 Ses.5 Ses.6 Ses.7 Ses.8 Ses.9 Ses.10 Avg $\uparrow$
1993 w/o CA 85.71 81.11 75.84 72.92 68.85 64.21 60.48 60.18 59.07 58.66 68.70
w/ CA 85.71 81.94 77.10 73.07 71.22 68.00 64.14 63.22 60.14 59.91 70.45
w/ SSCA 85.71 83.61 78.57 76.36 73.84 72.00 66.73 65.06 62.37 62.15 72.64
1996 w/o CA 81.99 77.34 71.36 64.74 66.14 63.83 62.54 60.25 60.09 58.39 66.67
w/ CA 81.99 78.06 74.57 67.91 67.15 64.57 64.59 62.62 61.92 62.08 68.55
w/ SSCA 81.99 78.06 74.81 69.78 71.18 67.9 66.43 65 64.2 64.19 70.35
1997 w/o CA 80.29 74.91 69.41 65.23 63.03 61.49 56.89 57.63 57.62 56.68 64.32
w/ CA 80.29 79.09 72 70.22 66.21 64.41 61.46 59.46 60.03 57.34 67.05
w/ SSCA 80.29 80.14 74.35 71.08 69.66 67.12 65.15 63.73 62.44 60.96 69.49
Table 17: Ablation results for unified classifier training and semantic shift estimation on ImageNetA. We report the overall performance of each session and the average performance.
Seed Method Ses.1 Ses.2 Ses.3 Ses.4 Ses.5 Ses.6 Ses.7 Ses.8 Ses.9 Ses.10 Avg $\uparrow$
1993 w/o CA 99.19 95.37 91.37 87.81 85.86 82.49 82.15 80.96 79.09 78.63 86.29
w/ CA 99.19 98.24 93.45 91.62 91.33 88.42 87.69 86.78 86.37 85.50 90.86
w/ SSCA 99.19 98.24 94.64 93.14 93.10 91.45 90.68 90.59 89.71 88.80 92.95
1996 w/o CA 100.00 94.79 91.09 90.02 88.34 85.92 85.59 82.74 80.87 79.05 87.84
w/ CA 100.00 95.83 94.60 92.60 92.00 90.83 90.56 86.84 85.40 85.58 91.42
w/ SSCA 100.00 96.25 96.06 94.96 94.30 93.81 93.32 91.26 90.22 89.10 93.93
1997 w/o CA 96.55 97.20 91.9 87.87 84.92 84.28 82.15 82.1 82.69 78.92 86.86
w/ CA 96.55 97.90 95.8 90.67 90.12 90.29 87.96 87.43 86.98 85.88 90.96
w/ SSCA 96.55 97.67 95.8 92.02 91.25 91.30 89.56 89.5 89.39 88.34 92.14
Table 18: Ablation results for unified classifier training and semantic shift estimation on CUB200. We report the overall performance of each session and the average performance.
Seed Method Ses.1 Ses.2 Ses.3 Ses.4 Ses.5 Ses.6 Ses.7 Ses.8 Ses.9 Ses.10 Avg $\uparrow$
1993 w/o CA 91.73 88.64 86.14 84.59 82.08 81.32 80.68 80.52 79.67 78.47 83.38
w/ CA 91.73 88.12 86.20 84.51 82.08 81.26 81.32 81.00 79.38 78.02 83.36
w/ SSCA 91.73 88.57 86.72 84.95 82.91 82.19 82.32 81.63 80.43 79.58 84.10
1996 w/o CA 89.55 86.63 85.27 82.45 81.64 80.36 79.56 78.33 78.34 77.40 81.95
w/ CA 89.55 87.59 86.37 83.58 82.15 81.12 79.85 78.81 78.51 77.67 82.52
w/ SSCA 89.55 87.11 86.65 84.04 83.34 82.00 81.51 79.87 79.73 78.72 83.25
1997 w/o CA 91.20 87.90 85.17 82.96 80.57 80.59 79.35 78.14 77.85 77.68 82.14
w/ CA 91.20 88.73 85.52 83.47 81.88 81.02 79.70 79.06 78.42 78.40 82.74
w/ SSCA 91.20 89.00 86.16 83.68 82.40 81.75 80.97 80.63 79.73 79.85 83.54
Table 19: Ablation results for unified classifier training and semantic shift estimation on ImageNetR. We report the overall performance of each session and the average performance.
C. More Implementation Details
To mitigate the impact of randomness in the experiments, we selected three different seeds (1993,1996, and 1997) to conduct experiments separately and calculate the average and variance. In experiments involving different PET methods, we fine-tuned the parameters inserted into the network without unified classifier retraining. For experiments of adapter dimension and layers insert, we conduct experiments on the ImageNetR dataset and we set the loss margin to $0.0$ and the scale to $20$. In the analysis of margin experiments, we set the scale to $20$ for all datasets. In the analysis of scale experiments, we set the scale to $0.0$ for all datasets.

### v2

@@ 5.4 Ablation Study
Baselines with different PET methods: Tab. VI shows the results of baselines with three different parameter-efficient tuning methods in each incremental session. It can be observed that the pre-trained model with an adapter achieves the best performance in terms of both the last session accuracy and average accuracy. Fig. 1 demonstrates that tuning with an adapter achieves a better balance between learning new classes and retaining knowledge of old classes. Both VPT-deep and SSF methods tend to prioritize learning new categories, which leads to increased forgetting of previously learned categories. Although VPT-shallow performs well on CIFAR, its limited parameters hinder the model from incrementally learning new classes on ImageNetR. More results on the other datasets can be found in the Supp.
@@
Unified classifier retraining vs. Separate local classifier: As we train separate task-specific classifiers in each incremental session, we propose to retrain the classifier to find the optimal decision boundary for all the classes. Tab. VII displays the ablation experiments of the classifier re-trained on ImageNetA which is the most difficult benchmark. It can be observed that whether it is a linear or a cosine classifier, retraining the classifier leads to a significant performance improvement. Additionally, incorporating the computation of prototype semantic shifts further enhances the performance by an additional $2\%$ in the cosine classifier. Compared to the classifier alignment methods that do not involve computing updated prototypes, our method demonstrates its superiority as the incremental stages progress. More results on the other datasets can be found in the Supp.
@@
Progressively tuning vs. first session adaptation: Tab. VIII shows the linear probing results of different adaption ways. After finishing the training of the last session, we freeze the pre-trained backbone and only train the classifier using all the samples. It is evident that not performing tuning and solely freezing the pre-trained model leads to the worst performance, regardless of the dataset. First-session adaptation proves to be a good choice as it reduces training time and works well for datasets like CIFAR100 and CUB200. However, for datasets such as ImageNetA and ImageNetR, which have significant domain gaps from the pre-trained model, relying solely on first-session adaptation is suboptimal. By continuously fine-tuning the adapter, we observe that the backbone exhibits stronger representability compared to only tuning in the first session.
@@
Different structures of the adapter: In this paper, we follow AdaptFormer [22] to use parallel adapterMLP as the adapter structure. We also delve deeper into different adapter structures such as Adapter [51] and Convpass [85]. Although these different tuning structures may exhibit performance differences under static settings, the performance differences among those adapter structures are minimal in the context of CIL shown in Tab. IX. This offers us the flexibility to employ various adapter structures within the context of the CIL paradigm.
Runtime cost of semantic shift. Performing semantic shift at each incremental session incurs extra computation time, and as the number of classes grows, the time cost gradually increases. Therefore, it is necessary to discuss the time costs associated with the semantic shift computation. We report the results of the time cost of the prototype-based method and sample-based method in each session shown in Fig. 7. It can be seen from the right figure that as the number of sessions increases, the time cost associated with the semantic shift continues to grow. Besides, it can be seen that the sample-based semantic shift method incurs significantly higher time costs compared to the prototype-based method across various datasets.
TABLE X: Comparison to traditional CIL methods on ImageNetR and ImageNetA dataset.
@@
SSIAT (Ours) $79.61_{\pm 0.40}$ $83.71_{\pm 0.42}$ $62.73_{\pm 1.01}$ $71.05_{\pm 1.73}$
Figure 8: The confusion matrix of our proposed method on session 2,4,7,9 on ImageNetA, CUB and ImageNetR datasets. The bottom-right quadrant of the intersection points of the decision boundaries depicts the novel classes being learned in the current session.
5.5 More Analysis
@@
Analysis of margin and scale: We conduct experiments on the ImageNetA dataset on the hyper-parameter of the scale and margin in our cosine loss as shown in Fig. 9 left. We discover that appropriately increasing scale can enhance the performance of the model when the margin is fixed. For example, when the margin is fixed at $20$, the performance is observed to continuously improve as the scale parameter is varied from $0.0$ to $0.5$. when the scale is set to $0.5$, the average accuracy is $3.23\%$ higher than when the value is $0.0$. We also conduct experiments to analyze the influence of the margin. As shown in Fig. 9 left. It can be observed that with the scale parameter held fixed, appropriately increasing the margin can also enhance the performance of the model. we also observe that different datasets have different appropriate margin values. More experimental results are presented in the Supp.
Adapter dimension and layers insertion: As shown in Fig. 9 right, we further conduct experiments on the specific position to insert the adapter module on the ImageNetA dataset. We observe that the performance is progressively improved with the number of layers increased when the middle hidden layer dimension is fixed. Thus, we insert adapter modules in all $12$ layers in our comparative experiments of various methods. We also conduct ablation experiments on the middle dimension in the adapter also in Fig. 9 right. We observe that increasing the dimension has a positive effect on the performance of the model when the position to insert adapter module is fixed. Interestingly, setting the middle dimension to 32 did not result in a significant decrease in performance. On the other hand, setting it to 256 led to an improvement in performance but also quadrupled the number of parameters. To strike a balance between performance and the number of fine-tuning parameters, we set the middle dimension to $64$.
Figure 9: The results of some hyperparameters. The left shows the effect of different combinations of margin and scale on the ImageNetA dataset. The right demonstrates the impact of varying the number of inserted adapter layers and the dimensionality of the adapters on performance.
TABLE XI: The CIL results of different semantic shift methods on ImageNetA.
[…]
CA[18] All 85.14 82.50 77.52 73.35 71.58 68.72 64.50 64.26 61.65 60.96 71.01
Instance-Based $l_{1}$ Distance [34] All 85.14 82.50 77.31 77.22 75.51 72.10 65.74 66.11 64.52 62.41 72.85
$l_{2}$ Distance [34] All 85.14 82.50 77.31 77.22 75.51 72.10 65.74 66.19 64.52 62.41 72.86
RoP-ILAP[86] k=100 85.14 82.50 77.31 70.92 66.35 60.51 53.43 49.84 46.95 44.63 63.75
RoP-ILAP[86] k=200 85.14 83.06 77.10 72.06 68.61 65.64 58.97 56.57 53.12 50.63 67.09
RoP-ILAP[86] k=300 85.14 83.06 77.73 73.93 71.22 67.79 61.20 60.98 57.28 54.84 69.31
RoP-ILAP[86] All 85.14 82.78 77.10 77.22 75.51 71.90 66.01 66.67 64.01 62.54 72.88
Prototype-Based PRAKA[87] All 85.14 80.56 73.74 66.48 65.16 60.82 57.36 54.57 52.26 50.56 64.66
SSIAT(Ours) All 85.14 82.78 77.31 77.51 75.03 71.79 66.10 66.59 64.23 63.00 72.94
TABLE XII: Results for different PET methods on ImageNetR/A.
@@
Different pre-trained models. We experiment with pre-trained models (PTMs) with different generalization abilities on four datasets shown in Tab. XIII. We conduct experiments on three different versions of ViT PTMs and it can be observed that our method generalizes well to various PTMs. Furthermore, The large-based ViT model can get better performance on ImageNetR and ImageNetA datasets. However, on other datasets, such as CIFAR100, the large-based ViT models may not necessarily achieve better performance.
Different tuning methods with SSCA. We observe that the adapter module exhibits the best performance with continuously fine-tuning. Therefore, we combined the semantic shift correction with continuous adapter fine-tuning, which led to even better results. To further substantiate the superior performance of the adapter, we also conduct experiments by integrating the semantic shift correction with other PET methods for comparison. We incorporate classifier alignment with semantic shift estimation into SSF and prompt tuning shown in Tab. XII. It can be seen that both the performance of prompt-based and SSF tuning approaches show significant improvement. However, our proposed method still outperforms them by a large margin. The results further verify the effectiveness of our proposed method.
Visualization of decision boundaries and confusion matrixes. To visually observe the changes in class decision boundaries during the incremental process, we conduct feature visualization on the CIFAR100 dataset. The visualization results are shown in Fig. 10. It can be seen that the discriminability of the base session features is excellent. After undergoing incremental training to accommodate the newly added classes, the discriminability of the new classes is also preserved, without affecting the discriminability of the base session classes. We show the confusion matrix on ImageNetA, CUB, and ImageNetR datasets in Fig. 8. We show the results across four sessions for each dataset. It can be seen that regardless of whether the current session involves learning new classes or previously learned classes, the model does not appear to exhibit significant confusion in its predictions.
TABLE XIII: Results on different pre-trained models on ImageNetR/A.
Pre-trained Model ImageNetR ImageNetA
ViT-base 1K $80.02_{\pm 0.30}$ $84.45_{\pm 0.42}$ $61.40_{\pm 0.78}$ $71.05_{\pm 1.20}$
ViT-base 21k $81.96_{\pm 2.60}$ $85.64_{\pm 1.24}$ $60.78_{\pm 1.52}$ $70.00_{\pm 1.06}$
ViT-large 21k $80.02_{\pm 0.30}$ $84.12_{\pm 0.21}$ $65.55_{\pm 3.76}$ $74.12_{\pm 3.16}$
Pre-trained Model CIFAR100 CUB200
ViT-base 1K $90.68_{\pm 0.23}$ $93.72_{\pm 1.13}$ $88.32_{\pm 0.48}$ $92.94_{\pm 0.55}$
ViT-base 21k $91.02_{\pm 0.87}$ $94.24_{\pm 0.88}$ $89.82_{\pm 0.59}$ $93.76_{\pm 0.58}$
ViT-large 21k $90.63_{\pm 0.69}$ $92.77_{\pm 1.96}$ $90.23_{\pm 0.05}$ $93.62_{\pm 0.14}$
Figure 10: The results of feature and decision boundary visualization on CIFAR100 dataset. The left shows five classes in the base session and the right shows ten classes after an incremental session.
Different methods of SSCA. There exist multiple methods available to calculate the similarity between features in SSCA. To compare the effectiveness of different SSCA approaches, we conduct experiments on ImageNetA dataset and the results are shown in Tab. XI. It can be seen that when incorporating the currently available data to correct semantic drift, various methods for computing similarity yield similar results. RoP-ILAP [86] utilizes the $k$-means algorithm to identify samples closest to the existing class prototypes and then calculates the corresponding semantic shift. We conduct experiments using the same method with varying values of $k$. The results indicate that as the value of $k$ increased, the average performance improved. This shows that incorporating all samples leads to better alleviation of semantic drift. Furthermore, it can be seen that our proposed prototype-based method yields better results compared to using the entire sample set. PRAKA employs an interpolation-based sampling technique to correct prototypes using features from the old class. However, the results indicate that its performance is moderate on imbalanced datasets such as ImageNetA.
6 Conclusion
Class-incremental learning on a pre-trained model has received significant attention and gained excellent performance in recent years. In this paper, we first revisit different PET methods in the context of CIL. Then, we propose that incrementally tuning the shared adapter and local classifier without constraints exhibits less forgetting and gains plasticity for learning new classes. Moreover, to train a unified classifier, we calculate the semantic shift of old prototypes and retrain the classifier using updated prototypes without access to past image samples in each session. The proposed method eliminates the need for constructing an adapter pool and avoids retaining any image samples. Experimental results on several CIL benchmarks and more general settings demonstrate the effectiveness of our method which achieves the SOTA performance.
Limitation. We demonstrate that the PET method offers significant advantages in CIL with pre-trained models, and adapter-based methods generally outperform other PET-based methods. However, further validation is needed for PET methods in real-world open-domain scenarios. Additionally, addressing semantic shift and retraining the classifier is required at each incremental session, leading to substantial time overhead as the number of sessions increases. Although our prototype-based semantic shift method reduces time overhead compared to other approaches, its application remains limited in high real-time requirements. Beyond ViT-based pre-trained models, research is increasingly focusing on vision-language multimodal pre-trained models for incremental learning tasks. Enhancing the performance of these multimodal models in more complex incremental learning settings will likely be a key area for future work.